Mr.Jon's garden covered 300 square feet. If the width was 5 feet less than the length, find the measurements of the garden.
Answer by ankor@dixie-net.com(22740)

Mr.Jon's garden covered 300 square feet. If the width was 5 feet less than the length, find the measurements of the garden.
:
Let x = length of the garden
then
(x-5) = width of the garden
:
Length * Width = 300
x*(x-5) = 300
:
x^2 - 5x = 300
:
x^2 - 5x - 300 = 0; our old friend, the quadratic equation
:
Fortunately, it will factor:
(x - 20)(x + 15) = 0
:
x = +20 is the length, (the positive solution is what we want here)
:
Width would be 15, obviously
:
Check our solutions: 20 * 15 = 300